An economic analyst employed by Agbesco Ltd observed that, 9.5 units of ‘Kalami’, Agbesco Ltd product was purchased by customers of Agbesco regardless of the price charged for the product. In addition,when Agbesco Ltd increased its price by GH1, quantity demanded of Kalami falls by 80%.

On the other hand,Agbesco Ltd supplied 80 units of Kalami irrespective of the price level. Interestingly, when Agbesco Ltd increased its price by GH1, quantity supplied of Kalami increased by 20%.

Write out an equation for the demand curve and supply curve for the company. Find the equilibrium price and quantity.

Sketch the demand and supply curves using their respective equation.

Calculate the price elasticity of demand and supply and interpret the results.

Calculate the producer, consumer and total surplus
1
Expert's answer
2019-11-11T15:34:50-0500

The equations for the demand curve and supply curve for the company are:

Qd = 95 - 0.8P,

Qs = 80 + 0.2P.

The equilibrium price and quantity are:

Qd = Qs,

95 - 0.8P = 80 + 0.2P,

P = 15,

Q = 83 units.

The price elasticity of demand and supply are:

Ed = -0.8×15/83 = - 0.14, so the demand is inelastic.

Es = 0.2×15/83 = 0.036, so the supply is inelastic too.

The producer surplus is:

PS = 15×(83 + 80)/2 = 1222.5.

The consumer surplus is:

CS = 0.5×(95 - 15)×83 = 3320.

The total surplus is TS = 1222.5 + 3320 = 4542.5.